OpenStudy (anonymous):

would the x intercept for f(x)=25^2+40x+16 be (-2/5,0)?

OpenStudy (anonymous):

scratch that...not the full equation

OpenStudy (anonymous):

i believe it should be (-4/5, 0), thats just with a glance though.

OpenStudy (anonymous):

joe

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

assuming the full equation should be: \[25x^2+40x+16\]

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

a^2+2ab+b^2

OpenStudy (anonymous):

(a+b)^2

OpenStudy (anonymous):

you can factor the equation into: \[(5x+4)^2 = 0 \iff 5x+4 = 0 \iff x = -\frac{4}{5}\]

OpenStudy (anonymous):

this same question has been posted again